OpenStudy (anonymous):

Will fan and medal! For the formula A^2+B^2=c^2 If I have sin (-sqrt3/3) What would be my a,b, and c in the end?

zepdrix (zepdrix):

Do you mean you have: \(\Large\rm\sin x=-\frac{\sqrt3}{3}\) Something like that?

OpenStudy (anonymous):

yes @zepdrix

OpenStudy (anonymous):

@freckles

zepdrix (zepdrix):

There are usually two clever ways to setup your side relations :) There are more options, but these are the most common way to do it. (Note: The hypotenuse should ALWAYS be positive). `Option 1` You can think of -sqrt3 as your opposite, and 3 as your hypotenuse. `Option 2` You can rewrite the entire fraction over 1 like this \(\Large\rm \dfrac{-\sqrt{3}/3}{1}\) and think of your opposite as -sqrt3/3 and your hypotenuse as 1.

zepdrix (zepdrix):

The first option is probably the most convenient though.

zepdrix (zepdrix):